### v1

@@ IV-A Framework Overview
6:        Compute loss: $\mathcal{L}=D_{\text{KL}}\left(\tilde{y}(x_{b})\,\Big{\|}\,\text{softmax}\left(\frac{\text{Distill\_Logits}(x_{b})}{T}\right)\right)$
7:        Update $M_{\text{distill}}$ using $SGD$ with $\eta$
8:     end for
9:     Update SWA model: $M_{\text{SWA}}\leftarrow\frac{n_{\text{SWA}}\cdot M_{\text{SWA}}+M_{\text{distill}}}{n_{\text{SWA}}+1}$
10:     $n_{\text{SWA}}\leftarrow n_{\text{SWA}}+1$
11:  end for
12:  Set $M_{\text{global}}^{(r+1)}\leftarrow M_{\text{SWA}}^{(r)}$
13:  return  Updated global model $M_{\text{global}}^{(r+1)}$
@@ V-A Datasets and Models
We conducted experiments on three widely used datasets: CIFAR-10, EMNIST, and Fashion-MNIST.
@@
Fashion-MNIST [23] comprises $70,000$ grayscale images of fashion products from $10$ categories, each of size $28\times 28$ pixels, providing a challenging alternative to MNIST dataset.
@@ V-C Attack Methods
TSBA [3] aims to evade anomaly detection mechanisms by carefully scaling the model weights after training. The adversarial client first trains its local model with the backdoor trigger until convergence. Then, it scales down the model updates by a factor $\epsilon$ that remains within acceptable bounds defined by the FL protocol. By doing so, the malicious updates appear similar to benign ones in magnitude, thereby bypassing defences that rely on detecting abnormal update sizes. The scaling factor $\epsilon$ is chosen to balance stealth and backdoor effectiveness.
@@ V-D Heterogeneous Setting
Extreme Heterogeneity: $\alpha=0.5$, $\alpha=0.3$, and $\alpha=0.1$. These lower values of $\alpha$ result in clients having data predominantly from a few classes, leading to highly Non-IID data distributions scenarios.
@@
By varying $\alpha$, we create a spectrum of data heterogeneity scenarios to comprehensively evaluate the robustness of the RKD framework under different degrees of Non-IID data distributions. This approach allows us to assess how well the RKD framework can handle challenges posed by data heterogeneity, which is critical for practical FL applications.
@@ V-E Evaluation Metrics
MTA measures the classification accuracy of the global model on a clean test dataset $\tau_{m}$, reflecting its ability to correctly predict the true labels of inputs without any backdoor influence. It is defined as: $\text{MTA}=\frac{\left\|\{x\in\tau_{m}\mid f(x)=y\}\right\|}{\left\|\tau_{m}\right\|},$ where $x$ is an input sample from the clean test dataset $\tau_{m}$. $f(x)$ denotes the prediction of the global model $f$ for input $x$. $y$ is the true label corresponding to $x$. ||$\cdot$|| ⋅ | represents the cardinality of the set.
@@
ASR evaluates the effectiveness of the backdoor attack by measuring the proportion of poisoned inputs that the global model misclassifies into the attacker’s target class. It is calculated on a test dataset containing backdoor triggers $\tau_{b}$: $\text{ASR}=\frac{\left\|\{x\in\tau_{b}\mid f(x)=y_{\text{target}}\}\right\|}{\left\|\tau_{b}\right\|},$ where $x$ is an input sample from the poisoned test dataset $\tau_{b}$ that contains backdoor triggers. $f(x)$ is the prediction of the global model $f$ for input $x$. $y_{\text{target}}$ is the target label that the attacker intends the model to output for backdoor inputs.
@@ V-F Experimental Results
V-F1 Defence Against A3FL Attack
Under highly heterogeneous non-IID conditions ($\alpha=0.3$), RKD demonstrated significant resilience against the A3FL attack on the CIFAR-10 and Fashion-MNIST datasets. As illustrated in Figures 1 and 2, RKD substantially reduced the ASR while maintaining high MTA compared to baseline methods.
@@ V-F1 Defence Against A3FL Attack
The superior performance of RKD against A3FL can be attributed to the synergy of its methodological components:
First, RKD employs cosine similarity to assess the directional alignment of model updates with the global model. RKD effectively detects malicious updates that embed backdoor triggers without significantly altering parameter norms—a limitation in methods like FLAME that cluster based on parameter magnitudes. This approach enables RKD to capture angular deviations introduced by adversarial models, isolating backdoor updates that might otherwise appear benign in terms of update size. Specifically, as detailed in our Automated Clustering methodology, the server computes the cosine similarity between each client’s model parameters and the global parameter vector. This metric captures significant directional deviations characteristic of malicious behaviour, allowing for effective detection.
Second, the use of HDBSCAN for clustering allows RKD to identify outliers without requiring a predefined number of clusters. By clustering the similarity scores obtained from cosine similarity, rather than the high-dimensional parameter vectors, RKD reduces computational complexity and avoids the curse of dimensionality. The dynamic adjustment of HDBSCAN’s minimum cluster size parameter, as described in our methodology, enables RKD to adapt to evolving model updates during training. This adaptability enhances clustering robustness, ensuring the accurate grouping of benign models while isolating potential malicious ones.
Third, after identifying and excluding malicious models through clustering, RKD utilizes the Model Selection component to further refine the aggregation process. By computing the median model parameters within the benign cluster and selecting models closest to the median model, RKD reduces the influence of any remaining outliers that may not have been filtered out during clustering. This selection is based on the distances of models from the median model parameters. As outlined in our methodology, this step mitigates the impact of potential undetected malicious updates and enhances the robustness of the aggregated model against backdoor attacks.
Fourth, the Knowledge Distillation Process plays a crucial role in consolidating the insights from the selected benign models. By distilling knowledge from the ensemble of these models into a global model using a temperature-scaled softmax function and the Kullback-Leibler divergence loss, RKD aligns the distilled model closely with the consensus of benign models. This process, as detailed in our methodology, mitigates the risk of incorporating hidden backdoor triggers that might not have been detected during clustering, ensuring that the global model retains high performance on legitimate tasks.
Integrating SWA during the distillation process further refines the global model by smoothing the training trajectory. SWA averages the weights of the model over multiple epochs, leading to convergence towards flatter minima in the loss landscape. This contributes to better generalization and strengthens the defence against backdoor attacks by reducing the model’s sensitivity to small perturbations introduced by adversarial updates. As per our methodology, SWA enhances the stability and generalization capability of the distilled model, ensuring robustness against backdoor attacks and maintaining high performance across Non-IID data distributions.
Crucially, RKD restricting access to the global model by disseminating the aggregated global model only to clients identified as benign. By withholding the updated global model from adversarial clients, RKD prevents attackers from adapting their strategies based on the latest model parameters. This measure is particularly effective against adaptive attacks like A3FL, which rely on feedback from the global model to refine their backdoor triggers. As described in our methodology, this approach limits the influence of malicious clients and enhances the overall security of the federated learning process.
By sending a minimally perturbed version of the global model (with a noise scale of $1\times 10^{-4}$) to suspected malicious clients, we obscured the exact state of the model, effectively preventing attackers from adapting their strategies. We evaluated the Perturbed Global Model (PGM) strategy within the RKD framework to maintain model performance while mitigating adaptive attack risks and reducing the likelihood of attackers realizing they have been identified. Comparing RKD with PGM (denoted RKD(PGM)) to RKD without PGM (denoted RKD) under various attack scenarios of A3FL, we found that RKD(PGM) maintained an average accuracy closely matching that of RKD, indicating that the minimal perturbations did not adversely affect performance.
@@ V-F3 Defense Against DBA Attack
V-F4 Defense Against TSBA Attack
The RKD framework robustly defends against TSBA on CIFAR-10 and EMNIST under Non-IID conditions ($\alpha=0.5$), as shown in Figures 7 and 8. RKD detects and mitigates TSBA manipulations, maintaining high accuracy and low ASR even with increased attacker ratios. Unlike other methods that falter under poisoned conditions, RKD excels with clean and poisoned datasets.
@@ V-F4 Defense Against TSBA Attack
The Impact of Heterogeneous Degree. We evaluated baseline defence methods and the RKD framework under varying Non-IID levels, including moderate ($\alpha=0.7$) and extreme ($\alpha=0.3,0.1$), as shown in Figure 9.
Baseline methods, despite high accuracy on clean inputs, falter in detecting embedded trigger patterns, leading to elevated ASR. In contrast, RKD consistently maintains robust defence, effectively countering the heightened threats of extreme heterogeneity. This highlights RKD’s superior adaptability and effectiveness where baseline defences struggle. While extreme heterogeneity impacts accuracy, RKD outperforms baseline methods by excluding backdoor models and ensuring robust performance and enhanced defence against backdoor attacks.
@@
Empirical Analysis of $Q$ Sensitivity
To evaluate the impact of the minimum cluster size ($Q$) on MTA and ASR, we performed experiments using the CIFAR-10 dataset under a Non-IID setting with $30$ clients, of which $40\%$ were malicious and executing A3FL backdoor attacks. The results are summarized in Table I.
TABLE I: Impact of $Q$ on MTA and ASR
$Q$ MTA ASR
Fixed $Q=2$ High High
Observations: Small cluster size allows malicious updates to dominate clusters, maintaining MTA but elevating ASR.
Fixed $Q=20$ Decreased, especially in later rounds Low
Observations: Larger clusters exclude many benign updates, reducing ASR. Over multiple rounds, this exclusion accumulates, leading to a global model that lacks sufficient diversity from benign updates, which negatively impacts the model’s ability to generalize and perform well on the main task.
Dynamic $Q$ Consistently high Low
Observations: Effectively excludes malicious updates while retaining benign ones, ensuring high MTA and reduced ASR.
Scalability. RKD enhances scalability by applying cosine similarity to model updates before clustering, transforming high-dimensional parameter vectors into scalar similarity scores. This dimensionality reduction significantly lowers computational complexity, making the clustering process more efficient. By avoiding clustering in the high-dimensional parameter space, RKD reduces both the time and resources required for defence operations. As shown in Table II, RKD’s defence time is $42.029$ seconds, substantially faster than FedDF and FedBE, which require $141.714$ and $198.765$ seconds, respectively. While RLR exhibits the shortest defence time, it compromises on detection accuracy due to its simplistic approach. FLAME is slightly more efficient in defence time, but RKD achieves a better balance between performance and robustness. These results highlight RKD’s overall efficiency and scalability, demonstrating that its methodological design—specifically the use of cosine similarity and efficient clustering—provides robust defence without incurring significant computational overhead.
@@
TABLE II: Comparative Defense Times
V-F5 Ablation Study
@@
RKD without Clustering and Model Selection. Removing both the Automated Clustering and Model Selection components, and leaving only the knowledge distillation process, significantly weakens the framework’s defenses. While it performs reasonably well under a $20\%$ DBA attack, as the proportion of adversarial clients rises to $40\%$, its defensive capabilities drop sharply. At a $60\%$ F3BA attacker ratio, the model suffers from severe performance degradation, with misclassifications aligning with attackers’ objectives. This highlights that without clustering, RKD is unable to effectively identify and isolate malicious updates, allowing backdoor attacks to compromise the global model.
RKD without Model Selection. Excluding the model selection component while retaining HDBSCAN clustering, this variant effectively manages a $20\%$ DBA attack ratio. However, with a $40\%$ attack ratio, performance is noticeably decreased. Under a $60\%$ F3BA attacker ratio, the removal of model selection further impairs RKD’s defense, resulting in increased ASR and reduced MTA. Even after the clustering phase effectively isolates most malicious updates, some malicious or anomalous updates may still be present. Without the median, these residual outliers can disproportionately influence the aggregated model, leading to instability in its parameters when evaluated under attack success rate. These findings underscore the critical role of model selection in refining the aggregation process by selecting models closest to the benign cluster centroid, thus enhancing robustness against higher proportions of attackers.
@@
V-G Additional Experiments
Defense Against A3FL Attack Under IID Conditions. As shown in Figure 11, RKD outperforms baseline models in defending against the A3FL attack on the CIFAR-10 dataset under IID conditions, across varying attacker ratios ($20\%$, $40\%$, and $60\%$). RKD’s robustness against adversarial attacks in IID settings underscores its effectiveness as a defence mechanism. The baseline models also demonstrate some inherent robustness at lower attacker ratios ($20\%$ and $40\%$), which can be attributed to the uniform data distribution in IID settings that allows models to learn more generalizable patterns, making it harder for attacks to succeed.
@@ V-G Additional Experiments
VI Conclusion and Future Work
Acknowledgments
-A Threat Model 1 - Malicious Clients Have Access to Benign Updates.
This scenario assumes an extremely informed adversary with capabilities far exceeding the typical assumptions in RKD threat models. Specifically, we consider the case where malicious clients have complete access to all benign updates in each round. While this represents an unrealistic adversarial scenario in federated learning, it serves as a valuable stress test to evaluate the robustness of RKD. Below, we outline the adversary coordination strategy: Adversary 1: Leveraging its knowledge of all benign updates, Adversary 1 selects its parameter update such that the new average, denoted as avg_new, is shifted to a specific point in the parameter space aligned with the adversarial objective. This manipulation ensures that avg_new reflects the desired direction of the attack while remaining plausible within the model update space. Adversary 2: Adversary 2 submits an update identical to avg_new. Since the clustering mechanism in RKD relies on cosine similarity values, Adversary 2’s update achieves a perfect cosine similarity score of 1 with avg_new, making it indistinguishable from benign updates.
Experimental Evidence.
Our experiments, illustrated in Figure 12, confirm that RKD’s reliance on cosine similarity clustering, while effective against most attacks, can be insufficient against such a highly coordinated strategy. Adversary 2’s inclusion in the benign cluster allows the adversaries to significantly influence the global model aggregation, causing a systematic shift in the global model’s parameters toward their objectives.
-B Threat Model 2 - Malicious Clients Do Not Have Access to Benign Updates.
This practical scenario assumes that malicious clients have only partial knowledge, meaning they lack direct access to benign updates and must approximate them indirectly. Adversary Strategy: In the Observation Phase, malicious clients behave benignly for multiple rounds, submitting updates that closely follow the global model to observe its evolution. In the Approximation Phase, they use these observations to estimate the average benign update, allowing them to generate updates that align with the attack strategy outlined in Threat Model 1.
Challenges for Adversarial Approximation:
Heterogeneity of Benign Updates. In FL, benign updates exhibit significant variability due to the non-IID nature of client data. This variability introduces inconsistencies that cannot be inferred from observing aggregated global updates alone. Additionally, the global model updates provide an oversimplified view of the collective contributions, masking the nuanced behavior of individual benign clients and limiting the accuracy of adversarial estimations.
RKD’s Clustering Based on Cosine Similarity. RKD clusters client updates using cosine similarity, emphasizing the directional alignment of each update with the global model. Even if adversaries attempt to approximate the average benign updates, achieving precise directional alignment across all malicious clients is inherently challenging. Malicious clients performing backdoor attacks typically have diverse and conflicting objectives, resulting in updates that do not uniformly align with the global model’s optimization direction. This variability makes it difficult for malicious clients to consistently approximate the benign update direction, preventing them from seamlessly blending into the benign clusters.
Moreover, HDBSCAN dynamically adjusts clustering boundaries based on the evolving distribution of similarity scores each round. This continuous adaptation makes it difficult for adversaries to predict the exact clustering outcomes.
RKD’s Knowledge Distillation Adds Randomness. RKD employs knowledge distillation using pseudo-labels generated by the ensemble on a validation dataset. This process introduces additional randomness into the final global model, disrupting adversarial efforts to align with benign updates.
Potential Impact on RKD.
The success of this attack hinges on the adversary’s ability to accurately estimate benign updates and replicate their alignment with the global model, thereby infiltrating benign clusters or evading detection.
Experimental Evidence.
Our experiments, illustrated in Figure 12, validate that RKD effectively defends against Threat Model 2. Despite malicious clients attempting to approximate benign updates, their approximations consistently deviate from actual benign updates. Even when the observation phase is prolonged, providing adversaries with additional iterations to gather data, their approximations remain inefficient and misaligned.
The accuracy trends of RKD and FedAvg under Threat Model 2, where adversaries lack direct access to benign updates and attempt to approximate them indirectly. This scenario simulates a more practical and realistic adversarial condition in federated learning. The results underscore the robustness of RKD and its superior performance in practical threat scenarios. At the same time, traditional aggregation methods like FedAvg in federated learning, lacking robust defense mechanisms, are unable to counteract even partially informed adversaries, resulting in a dramatic accuracy loss that underscores the necessity of designing effective defenses to ensure trustworthy global model aggregation in FL.
(a) Threat Model 1
(b) Threat Model 2
Figure 12: Impact of Threat Model 1 and 2 on Model Accuracy under RKD and FedAvg.
-C The Anti-Distillation Backdoor Attack (ADBA.)
The ADBA [26], initially designed for centralized machine learning scenarios utilizing knowledge distillation, has been adapted in this work to the federated learning paradigm. Originally, ADBA embeds a robust backdoor into a teacher model by dynamically optimizing a trigger and using a shadow model to ensure backdoor transfer through distillation. In extending this attack to FL, the global model is treated as the teacher, while client models act as distributed learners, effectively mirroring the hierarchical structure of knowledge distillation. Compromised clients inject backdoors by optimizing a crafted trigger using projected gradient descent while adhering to FL constraints. The results presented in 13 for RKD, in comparison with FedAvg and other baseline methods, underscore RKD’s effectiveness in detecting and mitigating ADBA backdoor attacks.
(a) $20\%$ of ADBA
(b) $40\%$ of ADBA
(c) $60\%$ of ADBA
Figure 13: Performance of baselines and RKD on CIFAR-10 under Non-IID settings against ADBA attacker clients.
-D Performance Under Non-IID and IID Conditions Without Attacks

### v2

@@ IV-A Framework Overview
7:        Compute loss: $\mathcal{L}=D_{\text{KL}}\left(\tilde{y}(x_{b})\,\Big{\|}\,\text{softmax}\left(\frac{\text{Distill\_Logits}(x_{b})}{T}\right)\right)$
8:        Update $M_{\text{distill}}$ using SGD with learning rate $\eta$
9:     end for
10:     Update SWA model: $M_{\text{SWA}}\leftarrow\frac{n_{\text{SWA}}\cdot M_{\text{SWA}}+M_{\text{distill}}}{n_{\text{SWA}}+1}$
11:     $n_{\text{SWA}}\leftarrow n_{\text{SWA}}+1$
12:  end for
13:  Set $M_{\text{global}}^{r+1}\leftarrow M_{\text{SWA}}$
14:  return  Updated global model $M_{\text{global}}^{r+1}$
@@ V-A Datasets and Models
We conducted experiments on three widely used datasets: CIFAR-10, EMNIST, and Fashion-MNIST:
@@
Fashion-MNIST [23] comprises $70,000$ grayscale images of fashion products from $10$ categories, each of size $28\times 28$ pixels, providing a more challenging alternative to MNIST dataset.
@@ V-C Attack Methods
The Anti-Distillation Backdoor Attack (ADBA)
The ADBA [26], initially designed for centralized machine learning scenarios utilizing knowledge distillation, has been adapted in this work to the federated learning paradigm. Originally, ADBA embeds a robust backdoor into a teacher model by dynamically optimizing a trigger and using a shadow model to ensure backdoor transfer through distillation. In extending this attack to FL, the global model is treated as the teacher, while client models act as distributed learners, effectively mirroring the hierarchical structure of knowledge distillation. Compromised clients inject backdoors by optimizing a crafted trigger using the PGD while adhering to FL constraints.
@@
TSBA [3] aims to evade anomaly detection mechanisms by carefully scaling the model weights after training. The adversarial client first trains its local model with the backdoor trigger until convergence. Then, it scales down the model updates by a factor $\gamma$ that remains within acceptable bounds defined by the FL protocol. By doing so, the malicious updates appear similar in magnitude to benign ones, thereby bypassing defenses that rely on detecting abnormal update sizes. The scaling factor $\gamma$ is chosen to balance stealth and backdoor effectiveness.
@@ V-D Heterogeneous Setting
Extreme Heterogeneity: $\alpha=0.5$, $\alpha=0.3$, and $\alpha=0.1$. These lower values of $\alpha$ result in clients having data predominantly from a few classes, leading to highly Non-IID data distribution scenarios.
@@
IID: In addition to the above Non-IID scenarios, we conduct experiments under IID conditions (i.e., when data is independently and identically distributed across clients) in the appendix to provide a baseline for comparison.
By varying $\alpha$ (and including IID experiments), we create a spectrum of data heterogeneity scenarios to comprehensively evaluate the robustness of the RKD framework under different degrees of Non-IID data distributions. This approach allows us to assess how well the RKD framework can handle the challenges posed by data heterogeneity, which is critical for practical federated learning applications.
@@ V-E Evaluation Metrics
[…]
@@
ASR evaluates the effectiveness of the backdoor attack by measuring the proportion of poisoned inputs that the global model misclassifies into the attacker’s target class. It is calculated on a test dataset containing backdoor triggers, denoted as $D_{\text{poison}}$: $\text{ASR}=\frac{\left\|\{x\in D_{\text{poison}}\mid f(x)=y_{\text{target}}\}\right\|}{\left\|D_{\text{poison}}\right\|},$ where $x$ is an input sample from the poisoned test dataset $D_{\text{poison}}$ that contains backdoor triggers, $f(x)$ is the prediction of the global model $f$ for input $x$, and $y_{\text{target}}$ is the target label that the attacker intends the model to output for backdoor inputs.
@@ V-F Experimental Results
V-F1 Defence Against A3FL Attack
Under highly heterogeneous Non-IID conditions ($\alpha=0.3$), RKD demonstrated significant resilience against the A3FL attack on the CIFAR-10 and Fashion-MNIST datasets. As shown in Figures 1 and 2, RKD achieved a substantially lower attack success rate while maintaining high accuracy compared to baseline methods.
These results confirm that RKD effectively distinguishes malicious from benign client updates and aggregates only reliable models. By restricting the dissemination of the updated global model to clients identified as benign, RKD prevents adversaries from adapting their strategies based on the latest global updates. In our primary approach—the Exclusion Strategy—malicious clients continue training with their current local models. Alternatively, the Perturbation Strategy (denoted as RKD (PGM)) provides suspected malicious clients with a minimally perturbed global model: $M_{\text{pert}}^{r+1}=M_{\text{global}}^{r+1}+\eta,$ where $\|\eta\|\approx 1\times 10^{-4}$. This slight perturbation effectively obscures the precise state of the global model, thereby limiting the opportunity for adaptive adversaries to refine their attacks.
Comparative experiments show that RKD (PGM) maintains an average accuracy nearly identical to that of RKD using the Exclusion Strategy, while still mitigating adaptive attack risks. Overall, the experimental findings illustrate that RKD robustly mitigates backdoor attacks under Non-IID conditions.
@@ V-F3 Defense Against DBA Attack
V-F4 Defense Against ADBA Attack
The RKD framework robustly defends against Anti-Distillation Backdoor Attacks (ADBA) on CIFAR-10 under Non-IID conditions ($\alpha=0.5$), as shown in Figure 7. Compared to FedAvg and other baseline methods, RKD effectively detects and mitigates ADBA backdoor attacks, demonstrating superior resilience and enhanced model integrity in challenging heterogeneous environments.
(a) $20\%$ of ADBA
(b) $40\%$ of ADBA
(c) $60\%$ of ADBA
Figure 7: Performance of baselines and RKD on CIFAR-10 under Non-IID ($\alpha=0.5$), evaluated against $20\%$, $40\%$, and $60\%$ ADBA attacker clients.
V-F5 Defense Against TSBA Attack
The RKD framework robustly defends against TSBA on CIFAR-10 and EMNIST under Non-IID conditions ($\alpha=0.5$), as shown in Figures 8 and 9. RKD detects and mitigates TSBA manipulations, maintaining high accuracy and low ASR even with increased attacker ratios. Unlike other methods that falter under poisoned conditions, RKD excels with clean and poisoned datasets.
@@ V-F5 Defense Against TSBA Attack
V-F6 The Impact of Heterogeneous Degree
We evaluated baseline defence methods and the RKD framework under varying degrees of data heterogeneity, including moderate ($\alpha=0.7$) and extreme ($\alpha=0.3,0.1$) Non-IID conditions (see Figure 10). Under extreme heterogeneity, many baseline methods achieve high accuracy on clean inputs but struggle to detect subtle backdoor triggers—resulting in elevated ASR. In contrast, RKD consistently maintains robust defence by effectively excluding malicious updates, which helps to suppress ASR while sustaining high MTA.
Notably, although extreme heterogeneity adversely impacts overall accuracy for all methods, RKD outperforms baseline defences by achieving a better balance between low ASR and high MTA. This indicates that a key contribution of our work is enhancing robustness in highly Non-IID scenarios. Moreover, under extremely Non-IID conditions ($\alpha=0.1$), baseline methods often struggle to generalize, resulting in model collapse that leads to a low ASR—since their offline behavior prevents an accurate assessment of robustness. In contrast, RKD sustains stable learning and robust defence, achieving both high accuracy and a genuinely low ASR.
Overall, these results highlight RKD’s superior effectiveness in challenging heterogeneous data environments.
@@ V-F6 The Impact of Heterogeneous Degree
V-F7 Empirical Analysis of $Q$ Sensitivity
We evaluated the impact of the minimum cluster size $Q$ on Main Task Accuracy (MTA) and Attack Success Rate (ASR) using the CIFAR-10 dataset under a Non-IID setting with 30 clients, 40% of which were malicious and executing A3FL backdoor attacks. Figure 11 presents the results.
Figure 11: Impact of $Q$ on MTA and ASR.
When $Q$ is fixed at 2, the resulting small clusters allow malicious updates to dominate, leading to a high ASR despite a relatively high MTA. In contrast, fixing $Q$ at 20 excludes many benign updates, yielding slightly lower accuracy compared to the settings with $Q=2$ or a dynamically adjusted $Q$, while still exhibiting a higher ASR due to the misclassification of some malicious updates.
A dynamic adjustment of $Q$ effectively mitigates these issues by excluding malicious updates while retaining the majority of benign ones, thus ensuring consistently high MTA and low ASR. These findings underscore the importance of dynamically tuning $Q$ to reduce the influence of residual outliers and adversarial updates, thereby preserving the overall robustness and performance of the global model.
V-F8 Scalability Analysis
RKD enhances scalability by applying cosine similarity to model updates before clustering, transforming high-dimensional parameter vectors into scalar similarity scores. This dimensionality reduction significantly lowers computational complexity, making the clustering process more efficient. By avoiding clustering in the high-dimensional parameter space, RKD reduces both the time and resources required for defence operations. As shown in Table I, RKD’s defence time is $42.029$ seconds, substantially faster than FedDF and FedBE, which require $141.714$ and $198.765$ seconds, respectively. While RLR exhibits the shortest defence time, it compromises on detection accuracy due to its simplistic approach. FLAME is slightly more efficient in defence time, but RKD achieves a better balance between performance and robustness. These results highlight RKD’s overall efficiency and scalability, demonstrating that its methodological design—specifically the use of cosine similarity and efficient clustering—provides robust defence without incurring significant computational overhead.
@@ V-F8 Scalability Analysis
TABLE I: Comparative Defense Times
V-G Ablation Study
@@
RKD without Clustering and Model Selection (Median). Removing both the Automated Clustering and Model Selection components, and leaving only the knowledge distillation process, significantly weakens the framework’s defenses. While it performs reasonably well under a $20\%$ DBA attack, as the proportion of adversarial clients rises to $40\%$, its defensive capabilities drop sharply. At a $60\%$ F3BA attacker ratio, the model suffers from severe performance degradation, with misclassifications aligning with attackers’ objectives. This highlights that without clustering, RKD is unable to effectively identify and isolate malicious updates, allowing backdoor attacks to compromise the global model.
RKD without Model Selection (Median). Excluding the model selection component while retaining HDBSCAN clustering, this variant effectively manages a $20\%$ DBA attack ratio. However, with a $40\%$ attack ratio, performance is noticeably decreased. Under a $60\%$ F3BA attacker ratio, the removal of model selection further impairs RKD’s defense, resulting in increased ASR and reduced MTA. Even after the clustering phase effectively isolates most malicious updates, some malicious or anomalous updates may still be present. Without employing the median to mitigate the influence of residual outliers, these outliers can disproportionately affect the aggregated model, resulting in unstable ASR measurements. These findings underscore the critical role of model selection in refining the aggregation process by selecting models closest to the benign cluster centroid, thus enhancing robustness against higher proportions of attackers.
@@
VI Conclusion and Future Work
Acknowledgments
Appendix A Additional Experiments under IID
A-A Defense Against A3FL Attack Under IID Conditions.
As shown in Figure 13, RKD outperforms baseline models in defending against the A3FL attack on the CIFAR-10 dataset under IID conditions, across varying attacker ratios ($20\%$, $40\%$, and $60\%$). RKD’s robustness against adversarial attacks in IID settings underscores its effectiveness as a defence mechanism. The baseline models also demonstrate some inherent robustness at lower attacker ratios ($20\%$ and $40\%$), which can be attributed to the uniform data distribution in IID settings that allows models to learn more generalizable patterns, making it harder for attacks to succeed.
@@ Appendix A Additional Experiments under IID
A-B Defense Against F3BA Attack Under IID Conditions.
(a) $20\%$ of F3BA
(b) $40\%$ of F3BA
(c) $60\%$ of F3BA
A-C Performance Under Non-IID and IID Conditions Without Attacks